# Electricity Market Bidding for Renewable Electrolyzer Plants: An Opportunity Cost Approach

### v2

@@ Abstract
Hydrogen produced through electrolysis with renewable power is considered key to decarbonize several hard-to-electrify sectors. This work proposes a novel approach to model the active electricity market participation of co-located renewable energy and electrolyzer plants, based on opportunity-cost bidding. While a renewable energy plant typically has zero marginal cost, selling power to the grid carries a potential opportunity-cost of not producing hydrogen when it is co-located with a hydrogen electrolyzer. We first consider only the electrolyzer, and derive its revenue of consuming electricity based on the non-convex hydrogen production curve. We then consider the available renewable energy production and form a piece-wise linear cost curve representing the opportunity cost of selling (or revenue from consuming) various levels of electricity. This cost curve can be used to model a stand-alone electrolyzer or a co-located hydrogen and renewable energy plant participating in an electricity market. Our case study analyzes the effects of market-bidding electrolyzers on electricity markets and grid operations. We compare two strategies for a co-located electrolyzer-wind plant; one based on the proposed bid curve and one with a more conventional fixed electrolyzer consumption. The results show that electrolyzers that actively participate in the electricity market lower the average cost of electricity and the amount of curtailed renewable energy in the system compared with a fixed consumption case. However, the difference in total system emissions between the two strategies is insignificant. The specific impacts vary based on electrolyzer capacity and hydrogen price, which determines the location of the co-located plant in the electricity market merit order.
@@ I Introduction
Although hydrogen produced through electrolysis enables low-carbon fuels and chemicals, the build-out of electrolyzers will cause significant increases in power demand [4]. Theoretically, electrolyzers can be flexible and controllable loads, and could fit well in modern power systems with high penetrations RES and limited transmission capacities [5]. In particular, if electrolyzers align their electricity consumption with the availability of renewable energy, their power demand might contribute to lift power prices and reduce the curtailment of renewable power, supporting the integration of renewable generation and improving the business case for continued expansion of renewable generation resources. However, while electrolyzers could theoretically be flexible, it is unclear if existing technology and economic incentives can encourage them to use this flexibility in a way that supports renewable generation. To mitigate these potential downsides, both the E.U. and U.S. have incentives that seek to tie hydrogen production to renewable generation [6, 7]. The primary goal of electrolyzers is to maximize profits from hydrogen production, creating a strong incentive to operate at high utilization rates to amortize the high capital costs associated with their construction. Further, electrolyzers might operate with Power Purchase Agreements (PPAs) and utility contracts that reduce their sensitivity to the real-time electricity price. Electrolyzers might therefore act more like large, constant loads than flexible resources, which could increase the cost of generation and emissions from the power system.
This work seeks to explore the potential benefits of active electricity market participation by electrolyzers, considering the viewpoint of both the electrolyzers and the grid. We derive cost curves that can model the bidding behavior of electrolyzers in electricity markets, specifically considering a Renewable Electrolyzer Plant (REP) consisting of a co-located RES and electrolyzer with a common grid connection. Since co-location setup reduces the dependency on the transmission grid and allows for direct consumption of renewable power, it is an economically and operationally attractive configuration for large-scale electrolyzer plants. We then analyze the impact of active market participation of large-scale electrolyzers on electrolyzer profits and power system operations.
@@ I-A Literature review
The market-based scheduling of electrolyzers or REPs is typically modeled through a profit-maximization problem, assuming a day-ahead power price forecast. References [8], [9], [10], and [11] consider different modeling approaches for electrolyzer or REP scheduling, with the focus of accurately representing the physics of the electrolyzer. These papers point to the importance of accurately representing the non-convex hydrogen production curve, which defines the hydrogen produced as a function of the electricity consumed. They perform optimal scheduling under the assumption of an exogenously determined (yet typically time-varying) electricity price, but do not account for the impact of electrolyzer scheduling on electricity network constraints and the electricity price formation. As the electrolyzer capacity expands, the consumption of electrolyzers will increasingly affect grid congestion and cause the electricity price to change. As a result, assuming that the electricity price is an exogenous signal as in [8]-[11] can lead to infeasible or suboptimal schedules, misleading estimates of profits, and negative impacts on the power system operations.
@@
The impact of electrolyzers on power and energy systems has been explored primarily through capacity expansion models. A technology-agnostic study was conducted in [12], employing a generation capacity expansion model to evaluate the role of “demand-sinks”, including, e.g., electrolyzers, in the grid under various cost assumptions. The study finds that demand-sinks might enable additional renewable capacity in the power system while having little effect on the cost of electricity. The demand sinks are assumed to be flexible and are dispatched assuming a fixed utility per consumed energy. Reference [13] explores the impact of various hydrogen investment support policies on the interaction of hydrogen, electricity and emission markets via a partial equilibrium model. Papers [14, 15], and [16] investigate the emissions associated with grid-connected electrolytic hydrogen production, evaluating the effects of proposed regulations such as the temporal matching between electrolyzer consumption and renewable production and the additionality of matched renewable sources. These papers represent the transmission network as either a zonal (i.e., modeling transmission capacities between regions only) or copper-plate (i.e., unconstrained transmission capacities) system. However, [17] argues that not considering the deliverability of renewable energy, i.e., through a zonal or copper-plate representation of the transmission network, drastically changes the analysis of emissions associated with a flexible load, such as a grid-connected electrolyzer.
The literature lacks investigations of the effects of electrolyzers (or REPs) in the power system that accurately represent both the hydrogen production curve of the electrolyzer and models the intra-regional transmission constraints of the power system. Further, there has been limited analysis on the difference in effects when the electrolyzer is operated with a fixed power consumption versus when its implicit flexibility is leveraged through active market participation.
@@ I-B Novelty and contribution
To address the identified gaps in the literature, we make two main contributions. First, we derive the marginal cost curve of REPs (or stand-alone electrolyzers), capturing the non-constant efficiencies of the hydrogen production curve and the time-varying availability of renewable generation. The cost curve is then converted to a bid curve that can be used for active participation in day-ahead and real-time electricity markets, representing either the opportunity cost of selling renewable generation to the grid rather than consuming it for hydrogen production or the revenue derived from importing electricity (if there is insufficient local renewable generation available). Second, we demonstrate how the electrolyzer bid curve can be used to analyze the scheduling of REPs in electricity markets under perfect competition. Specifically, we compare a market-bidding REP, which bids based on the derived cost curve, with a fixed-consumption REP, where the electrolyzer operates at a constant load. Our analysis examines how these strategies affect generation costs, renewable power curtailment, and system emissions, highlighting the influence of the REP’s marginal cost and its placement within the overall system merit order. Finally, we investigate how transmission constraints impact the perceived effects of connecting an electrolyzer to the power system.
@@
(c) REP cost curve
(d) REP marginal cost curve
@@
Figure 1: Derivation of the REPs (marginal) cost curve from the empirical hydrogen production curve. Illustrated for $\bar{P}^{\rm h}=P^{\rm RES}=500$ MW.
II Methodology
@@
This section first presents the derivation of the REPs’ continuous, non-convex, (marginal) opportunity cost curve from an empirical hydrogen production curve. We then approximate this cost curve as a convex, piecewise linear cost curve suitable for electricity market bidding. Finally, we present the mathematical formulation of our assumed electricity market clearing model, a DC Optimal Power Flow (DC OPF) problem.
II-A Derivation of the REP’s Opportunity-Cost Curve
@@
To model the REP as an active electricity market participant, we first derive its cost curve. Typically, renewable generators have zero marginal cost, as there is no variable cost associated with production. However, in a REP that includes both renewable generation and an electrolyzer, the situation is different. While producing hydrogen, the REP derives a revenue from consuming electricity. Selling this electricity to the grid instead would thus be associated with an opportunity cost, leaving the REP with a non-zero cost of exporting electricity. In the following, we derive a cost curve for the REP based on the opportunity cost associated with lost hydrogen production. The cost curve can be used to model the REP’s bids for electricity market-clearing models. From the cost curve, we further derive the REPs marginal cost curve. The steps of the derivation are described in detail below and illustrated in Fig. 1.
II-A1 Hydrogen production curve

### v3

@@ Abstract
Hydrogen produced through electrolysis with renewable power is considered key to decarbonize several hard-to-electrify sectors. This work proposes a novel approach to model the active electricity market participation of co-located renewable energy and electrolyzer plants, based on opportunity-cost bidding. While a renewable energy plant typically has zero marginal cost, selling power to the grid carries a potential opportunity-cost of not producing hydrogen when it is co-located with a hydrogen electrolyzer. We first consider only the electrolyzer, and derive its revenue of consuming electricity based on the non-convex hydrogen production curve. We then consider the available renewable energy production and form a piecewise linear cost curve representing the opportunity cost of selling (or revenue from consuming) various levels of electricity. This cost curve can be used to model a stand-alone electrolyzer or a co-located hydrogen and renewable energy plant participating in an electricity market. Our case study analyzes the effects of market-bidding electrolyzers on electricity markets and grid operations. We compare two strategies for a co-located electrolyzer-wind plant; one based on the proposed bid curve and one with a more conventional fixed electrolyzer consumption. The results show that electrolyzers that actively participate in the electricity market lower the average cost of electricity and the amount of curtailed renewable energy in the system compared with a fixed consumption case. However, the difference in total system emissions between the two strategies is insignificant. The specific impacts vary based on electrolyzer capacity and hydrogen price, which determines the location of the co-located plant in the electricity market merit order.
@@ I Introduction
Although hydrogen produced through electrolysis enables low-carbon fuels and chemicals, the build-out of electrolyzers will cause significant increases in power demand [4]. Theoretically, electrolyzers can serve as flexible and controllable loads, making them well-suited for modern power systems characterized by high shares of RES and limited transmission capacity [5]. Demand-side flexibility and price elasticity are widely recognized as key tools for enhancing the efficiency of electricity markets [6]. They can also strengthen the business case for RES by mitigating low electricity prices and reducing renewable curtailment. Although electrolyzers are theoretically flexible, it remains unclear whether existing technologies and economic incentives are sufficient to encourage their use of this flexibility in ways that support renewable generation. The primary goal of electrolyzers is to maximize profits from hydrogen production, creating a strong incentive to operate at high utilization rates in order to amortize the substantial capital costs of their construction. Further, electrolyzers might operate with fixed-price Power Purchase Agreements (PPAs) and utility contracts that reduce their sensitivity to spikes in the wholesale electricity price [7]. Electrolyzers might therefore act more like large, constant loads than flexible resources, which could increase the cost of generation and emissions from the power system. To mitigate these potential drawbacks, particularly the risk of increased emissions, the E.U. and the U.S. have introduced incentive-based regulations aimed at linking hydrogen production to renewable energy generation [8, 9].
This work seeks to explore the potential benefits of active electricity market participation by electrolyzers, considering the viewpoint of both the electrolyzers and the grid. We derive cost curves that can model the price-taking bidding behavior of electrolyzers in electricity markets, specifically considering a Renewable Electrolyzer Plant (REP) consisting of co-located RES and an electrolyzer with a common grid connection, as illustrated in Fig. 1. Since the co-location setup reduces the dependency on the transmission grid and allows for direct consumption of renewable power, it is an economically and operationally attractive configuration for large-scale electrolyzer plants. We then examine how the active market participation of large-scale electrolyzers affects both their profitability and the operation of the power system.
Figure 1: The renewable electrolyzer plant, REP, consisting of co-located RES and an electrolyzer with a common connection point to the grid. The REP generates electricity from the RES and produces hydrogen via the electrolyzer. The electricity produced can be used internally for hydrogen production or exported to the grid. When RES production is low, the REP may import electricity to maintain hydrogen production. The hydrogen produced is exported to an external demand, which is not specified or further considered here.
@@ I-A Literature review
The market-based scheduling of electrolyzers or REPs is typically modeled as a profit-maximization problem, based on a forecast of day-ahead electricity prices. References [10][11][12], and [13] consider different modeling approaches for scheduling electrolyzers or REPs, with a particular focus on modeling the physical characteristics of the electrolyzer. These references highlight the importance of accurately representing the non-convex hydrogen production curve, which describes the relationship between electricity consumption and hydrogen output. They perform optimal scheduling under the assumption of exogenously determined, though typically time-varying, electricity prices, without accounting for the impact of electrolyzer operation on electricity network constraints and electricity price formation. As electrolyzer capacity expands, their electricity consumption will increasingly contribute to grid congestion and influence electricity prices. As a result, assuming that the electricity price is an exogenous signal, as in [10][13], can lead to infeasible or suboptimal schedules, misleading profit estimates, and negative impacts on power system operations. Reference [14] develops a bi-level optimization model for a hybrid RES-battery-electrolyzer system. This approach endogenizes market prices by removing the assumption of an exogenous price signal, while assuming that the hybrid system participates in the market as a strategic bidder.
@@
The impact of electrolyzers on power and energy systems has been explored primarily through capacity expansion models. A technology-agnostic study was conducted in [15] employing a generation capacity expansion model to evaluate the role of “demand-sinks”, including electrolyzers, in the grid under various cost assumptions. Demand sinks are assumed to be flexible and dispatched according to a fixed utility per consumed energy. References [16, 17], and [18] investigate the emissions associated with grid-connected electrolytic hydrogen production, evaluating the effects of proposed regulations such as temporal matching between electrolyzer consumption and renewable generation, as well as the additionality of matched renewable sources. Reference [19] explores the impact of various hydrogen investment support policies on the interactions among hydrogen, electricity, and emission markets using a partial equilibrium model. However, these studies generally neglect to accurately model the active market participation and price-sensitivity of electrolyzers, focusing instead on simplified assumptions such as fixed loads, which can lead to overlooking the system-wide impact of their operational flexibility. In addition, these references represent the transmission network either as a zonal system (modeling transmission capacities between regions only) or as a copper-plate system (assuming unconstrained transmission capacities). Yet, as argued in [20], ignoring the deliverability of renewable energy, by relying on a zonal or copper-plate representation of the transmission network, can significantly distort the analysis of emissions associated with flexible loads such as grid-connected electrolyzers.
Based on the above literature review, the following research gaps are identified. Existing studies that focus on accurate electrolyzer modeling typically rely on either exogenous price assumptions or strategic bidding behavior. However, there is a lack of techno-economic modeling frameworks enabling the accurate representation of REPs into standard market-clearing formulations while still adequately capturing their key physical characteristics. This gap limits the ability to realistically assess the system-wide impact of REPs. The literature examining the impact of electrolyzers (or REPs) on the power system typically adopts simplified electrolyzer models and overlooks intra-regional transmission constraints. Moreover, while several studies investigate the effects of proposed regulations for electrolytic hydrogen production on system emissions and costs, there is limited analysis comparing the system-level effects of electrolyzers operating under fixed power consumption versus leveraging their inherent flexibility through active market participation. By deriving the true marginal cost curve of the REP, its scheduling can be directly simulated within electricity market-clearing models, under the assumption of perfect competition, without relying on exogenously determined electricity prices. This cost curve represents the REP as an active market participant and fully captures its price elasticity.
Similar marginal cost curves, often represented as price-quantity bidding curves, have been developed in the literature for various assets, or portfolios of assets. For example, the cost curve of thermal generators captures varying marginal production costs [21, 22], whereas the cost curve of retailers reflects energy procurement costs from the day-ahead market along with balancing costs [23]. For generic portfolios of assets, this model represents opportunity costs arising from coordinated bidding across sequential markets under uncertainty [24]. One example is the price-quantity bid curve of price-taking wind power producers under uncertainty, accounting for asymmetric imbalance costs [25]. Our proposal follows a similar direction but is novel in its focus on price-taking REPs, whose bidding model in the form of price-quantity curves has not yet been addressed in the literature. Unlike previous works that derive bid curves for generic portfolios of assets, our approach explicitly targets the unique techno-economic characteristics of REPs. The proposed bidding curve simultaneously captures the joint physical constraints and flexibility of the co-located renewable and electrolyzer assets, while modeling the cross-sectoral coordination between the power and hydrogen systems, an aspect largely overlooked in existing studies.
@@ I-B Novelty and contribution
To address the identified gaps in the literature, we make two main contributions. Our first contribution is a practical modeling framework that derives and integrates the bidding models of REPs into market-clearing formulations, providing the basis to derive actionable insights into their system-wide impacts on emissions, costs, and congestion, under perfect competition. To do so, we analytically derive the marginal cost curve of REPs or stand-alone electrolyzers, capturing the non-constant efficiencies of the hydrogen production curve and the time-varying availability of renewable generation. The cost curve is then converted to a bid curve that can be used for active participation in, for example, day-ahead electricity markets, representing either the opportunity cost of selling renewable generation to the grid rather than consuming it for hydrogen production, or the revenue from hydrogen production derived by importing electricity when local renewable generation is insufficient. Our second contribution builds on this framework by demonstrating its application: using the derived bid curves to derive the market-based scheduling of REPs, we offer a structured method to evaluate their system-wide impacts. Specifically, we compare a market-bidding REP, which bids based on the derived cost curve, with a fixed-consumption REP, where the electrolyzer operates at a constant load. Our analysis examines how these two distinct strategies affect the total generation cost of the system, renewable power curtailment, and system emissions, highlighting the influence of the REP’s marginal cost and its position within the overall system merit order. Finally, we investigate how transmission constraints impact the perceived effects of connecting an electrolyzer to the power system.
II REP and Market Modeling
In this work, we consider an REP that interacts with a short-term electricity market, as illustrated in Fig. 1. The REP is connected to the grid through a common connection point shared by both the wind farm and the electrolyzer. From the grid’s perspective, the REP appears as a renewable generator with variable power availability. However, the presence of the electrolyzer introduces a non-zero opportunity cost of selling power to the grid, as the power could instead be used to produce hydrogen. Further, the REP may become a net importer of electricity when wind production falls short of the electrolyzer’s capacity. Thus, the REP is also comparable to technologies like pumped storage hydropower plants.
The REP trades two products, electric power and hydrogen gas. The electric power is traded through an electricity market, e.g., a day-ahead market. The REP participates in this market through submitting a bid curve, which we later derive. We assume that the hydrogen is traded through a bilateral contract at a fixed price, but we do not specify the kind of contract apart from the fixed hydrogen price. In the following, we outline the main assumptions underlying the derivation of the REP’s bid curve and subsequent analysis, and briefly discuss their implications.
Electricity Market Assumptions
We derive the REP’s cost curve based on three main assumptions about the electricity market. First, the electricity market is centrally cleared with uniform pricing, which may be either nodal or zonal. Second, the market operates under perfect competition. Third, the electricity market requires submitted bid curves to be piecewise linear and convex (i.e. the price must increase with increasing production). While other bid formats, such as block orders used in E.U. markets, are possible, this assumption reflects the primary bidding format employed in several U.S. and E.U. markets [26, 27].
The first two assumptions imply that the optimal bid curve for any participant, including the REP, is to bid their true marginal cost [28]. Our bid curve derivation shows that the REP’s true cost curve must reflect the opportunity cost of selling electricity to the market instead of using it internally for hydrogen production. If the electricity market were cleared using pay-as-bid pricing, for example, the suggested bid curve would not be suitable, as it would not enable the REP to fully recover its investment costs.
The assumption of perfect competition often does not hold in practice. The REP, which benefits from both low-cost renewable power production and a large controllable load, might possess market power. However, exploiting market power is prohibited in all E.U. electricity markets under REMIT [29]. We argue that, similar to other generation technologies, the REP’s true marginal cost can be audited by regulators or system operators as part of market power mitigation efforts, which supports the validity of the assumption of marginal cost bidding.
Modeling of Renewable Energy Variability and Hydrogen Demand
Our derivation of the REP’s cost curve does not consider a specified hydrogen demand (such as a minimum demand over a defined period), since incorporating this would require modeling multiple future time steps and forecasting electricity market-clearing prices. Furthermore, we assume that co-locating the renewable energy source and the electrolyzer within the REP provides sufficient hedging against renewable production variability. As a result, the REP is treated as a fully merchant entity without relying on additional PPAs. However, in our case study, we compare this fully merchant REP, modeled using the derived cost curve, with a REP operating under fixed electrolyzer consumption.
The derived bid curve and case study are both deterministic and do not explicitly model the stochastic nature of the REP or other RES in the system. However, we do account for the variability of RES both in the derivation of the REP bid curve and within the case study.
Considered Time Horizon
The short-term temporal horizon adopted in this study excludes considerations related to optimal investment decisions, such as the size and location of the REP and its long-term profitability. Moreover, potential long-term implications, such as the REP’s influence on future investments in generation and transmission capacity, are not captured within the current framework. In addition, the analysis is limited to a single trading floor and does not consider the balancing market, re-dispatches, other forward markets, or reserve markets.
Modeling of Hydrogen Price
Hydrogen is typically traded through bilateral agreements, such as hydrogen purchase agreements, where the parties negotiate the price [30]. Therefore, in this work, we assume that the hydrogen price is fixed and exogenously determined. The hydrogen prices used are based on reports of hydrogen production costs, assuming that electrolytic hydrogen competes with conventionally produced hydrogen. The cost of conventionally produced hydrogen via reforming is naturally linked to the price of natural gas. For reference, the International Energy Agency (IEA) reported hydrogen production costs ranging from \$1 to \$4 per kg in 2021, with the upper bound rising above \$8 per kg in 2022 due to increases in natural gas prices [31]. We select a lower range value of \$1.5 per kg as the base case, aligning with cost data in the later presented case study from 2019. Additionally, we investigate a scenario where the REP receives a “green premium” in the form of a subsidy or production tax credit per kg of hydrogen, based on the highest proposed level of around \$4.5 per kg [8, 9]. These two prices, i.e., the market rate price of \$1.5 per kg and the subsidized price of \$6 per kg, represent lower and upper bounds of competitive hydrogen prices, excluding impacts such as carbon taxes. While time-varying hydrogen prices would directly influence the REP’s cost curve, we keep the cost of conventional gas generators constant in the case study. Since the market rate hydrogen price correlates with the gas price, a constant hydrogen price is deemed sufficient. Nonetheless, the proposed REP’s cost curve framework allows for time-varying hydrogen prices at the same temporal resolution as the electricity market considered.
@@ II REP and Market Modeling
(c) REP’s cost curve
(d) REP’s marginal cost curve
@@
Figure 2: Derivation of the REP (marginal) cost curve from the empirical hydrogen production curve. Illustrated for $\bar{P}^{\rm h}=P^{\rm RES}=500$ MW.
III Methodology
@@
This section begins by outlining the electrolyzer modeling approach used in this work, leading to an empirical hydrogen production curve. Based on this, it derives the REP’s continuous, non-convex, (marginal) opportunity cost curve. This cost curve is then approximated as a convex, piecewise linear function suitable for electricity market bidding. Finally, the section formulates the electricity market-clearing mechanism as a DC Optimal Power Flow (DC OPF) problem.
III-A Modeling of electrolyzers
This work considers low-temperature electrolyzers, such as proton exchange membrane or alkaline electrolyzers, which are both mature technologies. The electrolyzer produces hydrogen at various efficiencies depending on its partial load set-point, with a peak efficiency typically around 20 to 30 percent of nominal capacity for low-temperature electrolyzers. We base our electrolyzer model on the hydrogen production curve developed in [10], which does not have a closed form expression. Here, we first express the hydrogen production as a function of power consumption using an empirical model based on the works of [32] and [33], resulting in a nonlinear curve. We then approximate it as a piecewise linear, concave function, which is shown to be accurate even for a small number of pieces. The approximation is exact at the linearization points; otherwise it mainly underestimates the original non-concave curve. The largest approximation error is found at the lowest electrolyzer load, the non-concave part of the curve. However, errors at the lower part of the curve have less impact on the scheduling results, as the hydrogen production is naturally the lowest at low loads. Therefore, the approximation error appears to be negligible.
We assume that low-temperature electrolyzers have ramp rates that allow them to freely change their set-point within the time duration of the considered electricity market, which in this case is one hour. Although alkaline electrolyzers are generally slower than proton exchange membrane electrolyzers, ramp rates reaching full load within a minute have been reported for both technologies [34, 35]. However, in the case of a full shut-down, alkaline electrolyzers in particular may have start-up times that exceed one hour, which our model does not capture. To avoid full shut-downs, the electrolyzer might instead be operated in a stand-by mode, where it consumes a small amount of electricity (around 2.5% of nominal capacity), from which fast start up is possible [10]. Our derivation neglects the cost this stand-by consumption would incur. Furthermore, we assume that the electrolyzer plant is operated as a modular system, which enables us to neglect the lower load limit of individual electrolyzer modules [36].
III-B Derivation of the REP’s opportunity-cost curve
@@
To model the REP as an active electricity market participant, we first derive its cost curve. Typically, renewable generators have zero marginal cost, as there is no variable cost associated with production. However, in a REP that includes both renewable generation and an electrolyzer, the situation is different. While producing hydrogen, the REP derives a revenue from consuming electricity. Selling this electricity to the grid instead would thus be associated with an opportunity cost, leaving the REP with a non-zero cost of exporting electricity. In the following, we derive a cost curve for the REP based on the opportunity cost associated with lost hydrogen production. The cost curve can be used to model the REP’s bids for electricity market-clearing models. From the cost curve, we further derive the REP’s marginal cost curve. The steps of the derivation are described in detail below and illustrated in Fig. 2.
III-B1 Hydrogen production curve
The starting point of the derivation is the hydrogen production curve of the electrolyzer,
${\rm h}(p^{\rm h}).$ (1)